OpenStudy (anonymous):

5(3 + 2) = 15 + 10 is an example of the associative property for addition

OpenStudy (anonymous):

true or false

OpenStudy (mathstudent55):

Here are some properies. Now you decide. Associative: a + (b + c) = (a + b) + c Commutative: a + b = b + a Distributive: a(b + c) = ab + ac Identity: a + 0 = a